A Sunday Newspaper costs $2.50. Jerry used a combination of nickels, dimes, and
quarters to pay for it. He used 11 fewer dimes than nickels. He used exactly 3
quarters. How many nickels and dimes did he use?

Found 2 solutions by Edwin McCravy, Alan3354:
Answer by Edwin McCravy(20056)

Instead of doing your homework for you, here is another one exactly like yours.

A magazine costs $3.30. Jerry used a combination of nickels, dimes, and quarters
to pay for it. He used 14 fewer dimes than nickels. He used exactly 5 quarters.
How many nickels and dimes did he use?
0.05N + 0.10D + 0.25Q = 3.30
                    D = N - 14
                    Q = 5

Substitute (N - 14) for D and (5) for Q in 0.05N + 0.10D + 0.25Q = 3.30

0.05N + 0.10(N - 14) + 0.25(5) = 3.30
   0.05N + 0.10N - 1.40 + 1.25 = 3.30
                  0.15N - 0.15 = 3.30
                         0.15N = 3.45
                             N = 3.45/0.15
                             N = 23

Subsitute (19) for N in D = N - 14

D = (23) - 14
D = 9

So the answer is 23 nickels and 9 dimes.

Now do your problem the exact same way.

Edwin

Answer by Alan3354(69443)

$2.50 for a newspaper? Is that realistic?